OpenStudy (anonymous):

Multiply (x – 2)(-5x2 – 3x – 4)

OpenStudy (anonymous):

Multiply each term in the second bracket, firstly by the first term in the first bracket, and then by the second term in the first bracket.

OpenStudy (anonymous):

Hint: (a+b)(c+d)=a(c+d) + b(c+d)= a*c+a*d+b*c+b*d

OpenStudy (anonymous):

does it help?